A uniform electric field of magnitude 3.50x104 N/C makes an angle of 47 ° with a plane surface of area 1.63x10-2 m2 What is the electric flux through this surface? N m2/C Submit Request Ans
Answer #1

Solution:

Given data:

Electric field of magnitude( E ) = 2.90×10^4 N/C

Angle ( Theta ) =  47^{0}

Plane surface of area ( A ) =  1.43×10^−2 m^{2}

Now, we have to find out the:

The electric flux through this surface :

We already know the formula for the  electric flux is,

Electric flux  Phi = EA COS (Theta)

  Phi = ( 2.90×10^4 N/C ) ( 1.43×10^−2 m^{2}) COS( 47^{0})

= 414.7 N.m^{2}/ C (0.68189 )

= 282.8247  N.m^{2}/ C

= 2.82 x 10   N.m^{2}/ C

herefore Electric flux  Phi  through this surface = 2.82 x 10   N.m^{2}/ C